X1 , X2 and X3 is a random sample of size 3 from a population with mean mu and
variance sigma^2. T1, T2 and T3 are the estimators to estimate mu, and are given by
T1 = X1 + X2 - X3; T2 = 2X1 + 3X3 - 4X2 and T3 = 1/3(lambda X1 + X2 + X3).
(i) Are T 1 and T2 unbiased ? Give reason.
(ii) Find the value of A, such that T3 is
unbiased.
(iii) Which is the best estimator ? State
giving reasons.
1
Expert's answer
2020-09-13T18:17:09-0400

"(i) \text{We will show that } T_1 \text{ and } T_2 \text{ are unbiased estimators}\\
\text{of the parameter } \mu \text{ using the definition of an unbiased estimator}.\\
E(T_1)=E(X_1+X_2-X_3)=\mu+\mu-\mu=\mu.\\
E(T_2)=E(2X_1+3X_3-4X_2)=2\mu+3\mu-4\mu=\mu.\\
\text{So } T_1 \text{ and } T_2 \text{ are unbiased estimators of } \mu.\\
(ii) E(T_3)=\mu.\\
\frac{1}{3}(\lambda\mu+\mu+\mu)=\mu.\\
\text{Then }\lambda=1.\\
(iii) D(T_1)=D(X_1+X_2-X_3)=\sigma^2+\sigma^2-\sigma^2=\sigma^2.\\
 D(T_2)=D(2X_1+3X_3-4X_2)=2\sigma^2+3\sigma^2-4\sigma^2=\sigma^2.\\
 D(T_3)=D\big(\frac{1}{3}(X_1+X_2+X_3)\big)=\frac{1}{9}(\sigma^2+\sigma^2+\sigma^2)=\frac{1}{3}\sigma^2.\\
\text{Since } T_3 \text{ has the lowest variance then } T_3\\
\text{is the best estimator of } \mu."